Ticket Number: HEA-NOC/20071210-8 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20071210 15:59 UTC Problem Start: 20071212 13:00 UTC
Ticket Closed: 20080107 13:59 UTC Problem End: 20071212 15:00 UTC
Site/line: DIT, UCD
We plan to move our IPv6 tunnel endpoint to a new router that has been
deployed in Citywest as part of the new IP equipment.
IPv6 traffic to DIT and UCD will be interrupted for the duration of the
20071210 15:55 DW Contacted affect clients, open ticket.
20071211 13:14 DW Investigating problem with ISIS between 7200 and CRS.
This work may be postponed.
20071211 13:38 DW MTU mismatch found and resolved by lancomms. Proceeding
20071211 14:04 DW Could not complete within maintenance window. Scaling
back and will reschedule.
20071212 10:14 DW Having checked with affected clients, rescheduling
windows for 13:00-14:00 on Wed 12th and Thu 13th December.
20071212 15:17 DW Problems encountered with IPv6 connectivity to UCD both
in the before and after states, but customer has indicated that this
should not be an obstacle to the transition. Now using the new router, and
awaiting confirmation from DIT that there are no further problems.
20080107 13:58 DW DIT responded on the day, and there have been no
problems observed since. Closing ticket.
Time to Fix:
Sporadic outages during a 2 hour period.
This ticket can be monitored at http://www.hea.net/tickets/20071210-8
HEAnet Limited. Registered in Ireland, No. 275301.